Suppose the current exchange rate for the Russian ruble is RUB 64.18. The expected exchange rate in three years is RUB 69.32. Assume that the anticipated rate is constant for both countries. What is the difference in the annual inflation rates for the United States and Russia over this period? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

Difference in inflation between Russia and US =(Future Exchange rate/Spot rate)^(1/n)-1 =(69.32/64.18)^(1/3)-1 =2.60%

If Difference of US and Russia is given it could be =-2.60%